Thursday morning 09:30 Dump of pPb fill.

Slides:



Advertisements
Similar presentations
Beam commissioning strategy Global machine checkout Essential 450 GeV commissioning System/beam commissioning Machine protection commissioning.
Advertisements

LHC progress Saturday 2 nd May 2015 Coordination: Mike Lamont, Wolfgang Hofle.
Wednesday 12 th September Pumping, removal of vacuum equipment from tunnel in the morning TOTEM & ALFA, RP movement tests: OK 11:50 pre-cycle started,
21/05/2012LHC 8:30 meeting Overview Week 20 Injectors, LHCb polarity, Beam-Beam Coordinators: Bernhard Holzer, Jan Uythoven.
17/09/2012LHC 8:30 meeting LHC Summary Week 37 Coordination: Jorg Wenninger & Jan Uythoven.
LHC Progress Friday 30 th October 2015 Coordination Week 44: Massimo Giovannozzi, Wolfgang Hofle, Jorg Wenninger.
06:00 – Ongoing injection problems on beam 2  07:42 – Start of injection investigations  11:11 – Injection problem fixed. Resteering of transfer line.
Ion Commissioning: Thursday & Friday
Saturday Coordinators weekend: Mike Lamont, Eva Barbara Holzer
Wednesday :06 Dump fill 3533: interlock BPMS in IP6
-9:00: Test IP transverse adjustment (CMS) and optics verification.
Saturday 24 March 2012.
Wednesday 21 March 2012.
Monday :00 Stable beams 9_6_6_6 13:40 Access.
Friday 30th September 06:00 Operator dump fill #2171, int. lumi about 75 pb-1 Previous operator dump from physics was on 14 September ! 07:00 Start of.
MD Report 24 June 2012 Machine coordinators: Barbara Holzer, Mike Lamont MD Coordinators: Ralph Assmann, Giulia Papotti, Frank Zimmermann MD#2 News & Plan.
Friday 07:00 Power glitch on the EDF network. Point 8 affected.
Tuesday 11th September LHCb dipole presently fault off
Saturday 30th June - Sunday: Rocky !
LHC Morning Meeting - G. Arduini
Thursday morning – optics correction
Monday – recovery from MKI flashover
Thursday 17th May 01:11 Fill #2633 dump in squeeze arc circulator load channel Ad1B 03:16 Fill #2644 Stable beams, lumi 5.2e33 cm-2s-1 04:38 Dumped by.
Lost at 12:05 to power converter (RTQX2.R1) problem
Note: squeeze in Atlas, CMS, LHCb and Alice on going
Wed :40 Lost S78 - QPS controller.
Friday 31st August & night
Wednesday 8th August & night
Friday Morning 3:44 – 15:04 stable beams Fill (dumped by operator)
Tuesday 20 March 2012.
Week 35 – Technical Stop and Restart
Summary of Week 26 Main aims: G. Arduini, B. Holzer, M. Lamont
MD#2 News & Plan Tue – Wed (19. – 20.6.)
Monday Morning – recovery from cryo.
Friday morning 10:30 Optics measurements and corrections.
LHC Morning Meeting - G. Arduini
Not the most productive 24 hours on record
Friday Morning 7:30 injection for fill 2908
Friday 16/3 08:30 – 10:15 Access for QPS and EPC (BI in point 4)
Wednesday 09 March …:09:00: beta* measurement (1.5 m) by K-modulation and with AC-dipole excitation. 1/17/2019 LHC 8:30 meeting.
BI cont. *** Summary Matching Monitor test *** We only had ~15minutes of inject and dump, but managed to see a profile with ~9e9 p (!)
Wednesday 9:08 Fill 3214 dumped by OP after 5:26 hours in stable beams
Saturday 7th May Sat – Sun night
Tuesday morning 09:30 Beam dump, SIS dump due to lost connections with FCG FECs. Coincides with starting a development SIS server to test new interlocks.
Friday 23rd March 08:00 Access
Sunday/Monday & :30 Stable beams. Fill #1851.
Aims for the week Recover from technical stop and precycle at 10A/s
Monday :20 beta beat corrections during squeeze for flat machine finished Afterwards fixing issues at injection: BPM rephrasing for B2 solved.
Monday 14:00 beam dump Access:
Wednesday 10:00 test of the un-squeeze to 90 m at 4 TeV.
Tuesday morning 08:30-09:00 Squeeze with probe.
Summary week 18 Intensity ramp up & production
Monday 8th December Morning used for RF work (IP z, injection), damper noise checks. Stable beams period with LHCb spectrometer OFF – lost beam after.
Thursday 26th July 13:21 Fill 2880 beam dumped
Monday morning 09:40 Dump fill 2838, integrated ~130 pb-1.
Sat morning Until 12:40: Loss maps at injection
Friday 14th September 03:15 Beta* at 1 km. Optics measurements, see yesterday. 07:00 Dump. 08:00 Refill with 3 nominal bunches for high beta. 11:15 Beams.
Monday :15 fill 3523 dumped by BPMs IP6
Saturday 12 March Morning: problem on RSD1.A12B2 – flow rate problem on the cooling water for the converter. 14:00: Loss map -500 Hz (positive energy offset)
Wednesday 16th May 03:05 beam dump, RF fault crowbar problem on module M1B2 06:09 Another crowbar at injection. Access for RF; problem Beam Imminent Warning.
Wednesday 23/2 Thanks CRYO!!!.
Thursday
Monday 26th September 7:30 Beam dump, QSP on RQX.R1, lost cryo in ITR1, 91 pb-1. Cryo recovery, needed about 1.5 h. Setting of TI2 collimators to 5 sigma.
Another Immortal Fill….
Monday 27th August & night
Wednesday morning / afternoon
Wednesday morning Waiting for PS – problem on RF cavity.
Saturday 28th 08:17 Fill #1812: Stable beams, peak luminosity ~1.1x1033 cm-2s-1 19:00 Go to adjust for RF-checks; Record fill lumi of 31.7 pb-1 for CMS.
Tuesday
Wednesday :35 : Beam lost because of trip of RCBXH3.L1
Presentation transcript:

Thursday morning 09:30 Dump of pPb fill. RF intervention on klystron line 2B1. 11:30 Preparing for injection. Problem with RF settings – fixed by G. Kruk. Changes in settings made since BP was created were not propagated. 12:30 Again problem with 2B1. Switch off and remove from configuration of used cavities. 13:00 Injecting. BPM phasing (for bunch by bunch). SPS no beam – HV cable issue. 15:00 Ramp with 3 probes / beam. 12/2/2018 LHC 8:30 meeting

Thursday afternoon 15:40 Beam dumped in unsqueeze to 90m. Again the energy limits on TCTs… Too tight for 90m, and the limits were not opened for high beta. It was ‘assumed’ that we would mask everything (unfortunately not communicated). But this is and should not be(come) common practice. The SBF can have glitches… Opened energy limits for TCTs at 4 TeV (R. Bruce). And TCPs to reduce scraping in ramp. But TCT interlocks back in next cycles… (but now masked). Too many other problems to investigate further. 17:30 Filling again 19:30 Ramp, then squeeze to 90 m. At 90 m the orbit FB suddenly oscillates on B1V and kicks out ½ of the beam. Not understood at all… 12/2/2018 LHC 8:30 meeting

Thursday afternoon In parallel the unsqueeze to 500 m does not load. Current limits on RTQX in 1 and 5. Limits were modified following a MP event with a trip of the triplet circuit. The limits were set based on currents used in OP (ECR) to +- 100 A (need -300 A for 1 km). The currents used for 500/1 km were overlooked in the analysis. A change of limits requires PCs to be OFF… Beam2 lost when AC-dipole key was turned without masking BIS input (while sitting at 90 m). 20:30 Dump and try again. PC limits updated to +- 400 A. 12/2/2018 LHC 8:30 meeting

Thursday getting late 22:00 Ramp again. De-squeeze to 90 m fine. Then at start of unsqueeze to 500m orbit FB instability, orbit oscillates. Managed to control orbit by switching off FB when diverging, then back ON, some convergence and so forth. Why did it never diverge before, even up to 1 km? While trying to sort out OFB and run it with a different optics (with Ralph also in remote), OFB crash – this means all RT settings lost. Tried to save, but end of game no. 3. 12/2/2018 LHC 8:30 meeting

Thursday getting later and later 01:00 Back at injection. RS+JW checking out OFB. Nice convergence with injection optics, unstable with 1km optics (not surprising). No real anomaly. Injection optics - OK 1km – unstable or slow convergence 12/2/2018 LHC 8:30 meeting

Unstable with injection optics So late… At 90m. This time the optics load worked. Not understood why the same commands to send the optics did not seem to have any effect the last times. Switch optics OK Unstable with injection optics 12/2/2018 LHC 8:30 meeting

So late… At 500m. Again optics load worked, contrary to last time. With 90m optics With 500m optics Switch optics 12/2/2018 LHC 8:30 meeting

Finally 03:15 At 1 km – no losses in de-squeeze. 12/2/2018 LHC 8:30 meeting

Finally measurements at 1 km 03:30 – 6:30 beta beat / optics measurements at 1 km After applying the corrections calculated from the data taken in June 2012 we performed the measurements. The beta-beat went down from ~30% to less then 10% after applying the corrections. 12/2/2018 LHC 8:30 meeting

Optics measurements at 1 km Results Off-momentum measurements were also done to check the dispersion. No big difference between the measurement and the model was found. A k-modulation measurements was also performed to measure the beta* at IP1 and IP5. To be analyzed offline. 12/2/2018 LHC 8:30 meeting

Collisions 07:00 Collisions at ALPHA, about 1 Hz 07:18 Dump and refill to go for small indiv to take some data with pots in on the next fill, during EDF intervention Sune Jakobsen 12/2/2018 LHC 8:30 meeting

Plan Today at 09:00 EDF intervention on SPS HV cable. Then (options): Potentially no beam from SPS for the morning. Go for another 1000 m fill during EDF intervention RF line 2B1 repair (klystron modulator). Then (options): pPb MD many p bunches against Pb. Q20 in SPS. LHCb polarity. Test fill with split collision BP and LHCb Pol-. One more test fill needed if split collision BP to be used after TS3 in ramp up. Physics. 12/2/2018 LHC 8:30 meeting